OpenStudy (anonymous):

A father lefta will of $45 between his two daugthers aged 8,5years and 16 years such that they will get equal amounts when each of them reach the age 20 years. The original amount of $45 was invesed at 10% per annum simple interest. How much did the elder daughter get at the time of the will?

OpenStudy (anonymous):

The father gives the older and younger daughters A and B amounts respectively. From the question, it is known that A + B = 45. The money after x years can be calculated by the formula: Q = (P)*1.1^x Where Q is money after interest and P is money before interest. The older daughter will have to wait 4 years and therefore will have $(A*1.1^4) when she turns 20 The younger daughter will have to wait 11.5 years and will have $(B*1.1^11) when she turns 20 Note that it isn't 1.1^(11.5) because interest is only applied at the end of the year. Because both daughters receive the same amount at the ages of 20: A*1.1^4 = B*1.1^11 A*1.1^4 = (45-A)*1.1^11 A*(1.1^4+1.1^11) = 45*1.1^11 A = 45*1.1^11/(1.1^4+1.1^11) = 29.73 B = 45-A = 15.27 So the older and younger daughters are originally given $29.73 and $15.27 respectively.